OpenStudy (anonymous):

Ismael is comparing cell phone plans before upgrading his phone. Ameri-Mobile offers a low activation fee, but a high monthly payment. Cell-U-Later offers a lower monthly rate, but the activation fee is higher. Create a possible algebraic expression for both Ameri-Mobile and Cell-U-Later that shows the amount paid after an unknown amount of months have passed. Justify how you created those expressions, and identify what each term and factor represents in terms of the cell phone plans. Please Help

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@paki @nincompoop

OpenStudy (igreen):

You could say that Ameri-Mobile charges a $10 activation fee, and $15 a month. You can also say that Cell-U-Later charges a $20 activation fee, and $10 a month.

OpenStudy (igreen):

So your equation for Ameri-Mobile will be: y = 10 + 15x And your equation for Cell-U-Later will be: y = 20 + 10x Where x = months.

OpenStudy (igreen):

I think that answers your whole question..lol.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @Lexis1615

OpenStudy (anonymous):

@iGreen i got more

OpenStudy (igreen):

Okay, just close this one and open a new one.

OpenStudy (anonymous):

@iGreen give me 5 minutes lol